%apsrev4-2.bst 2019-01-14 (MD) hand-edited version of apsrev4-1.bst
%Control: key (0)
%Control: author (8) initials jnrlst
%Control: editor formatted (1) identically to author
%Control: production of article title (0) allowed
%Control: page (0) single
%Control: year (1) truncated
%Control: production of eprint (0) enabled
\begin{thebibliography}{25}%
\makeatletter
\providecommand \@ifxundefined [1]{%
 \@ifx{#1\undefined}
}%
\providecommand \@ifnum [1]{%
 \ifnum #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \@ifx [1]{%
 \ifx #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \natexlab [1]{#1}%
\providecommand \enquote  [1]{``#1''}%
\providecommand \bibnamefont  [1]{#1}%
\providecommand \bibfnamefont [1]{#1}%
\providecommand \citenamefont [1]{#1}%
\providecommand \href@noop [0]{\@secondoftwo}%
\providecommand \href [0]{\begingroup \@sanitize@url \@href}%
\providecommand \@href[1]{\@@startlink{#1}\@@href}%
\providecommand \@@href[1]{\endgroup#1\@@endlink}%
\providecommand \@sanitize@url [0]{\catcode `\\12\catcode `\$12\catcode
  `\&12\catcode `\#12\catcode `\^12\catcode `\_12\catcode `\%12\relax}%
\providecommand \@@startlink[1]{}%
\providecommand \@@endlink[0]{}%
\providecommand \url  [0]{\begingroup\@sanitize@url \@url }%
\providecommand \@url [1]{\endgroup\@href {#1}{\urlprefix }}%
\providecommand \urlprefix  [0]{URL }%
\providecommand \Eprint [0]{\href }%
\providecommand \doibase [0]{https://doi.org/}%
\providecommand \selectlanguage [0]{\@gobble}%
\providecommand \bibinfo  [0]{\@secondoftwo}%
\providecommand \bibfield  [0]{\@secondoftwo}%
\providecommand \translation [1]{[#1]}%
\providecommand \BibitemOpen [0]{}%
\providecommand \bibitemStop [0]{}%
\providecommand \bibitemNoStop [0]{.\EOS\space}%
\providecommand \EOS [0]{\spacefactor3000\relax}%
\providecommand \BibitemShut  [1]{\csname bibitem#1\endcsname}%
\let\auto@bib@innerbib\@empty
%</preamble>
\bibitem [{\citenamefont {Liang}\ and\ \citenamefont
  {Wang}(2005{\natexlab{a}})}]{Liang:2004ph}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Z.-T.}\ \bibnamefont
  {Liang}}\ and\ \bibinfo {author} {\bibfnamefont {X.-N.}\ \bibnamefont
  {Wang}},\ }\bibfield  {title} {\bibinfo {title} {{Globally polarized
  quark-gluon plasma in non-central A+A collisions}},\ }\href
  {https://doi.org/10.1103/PhysRevLett.94.102301} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {94}},\
  \bibinfo {pages} {102301} (\bibinfo {year} {2005}{\natexlab{a}})},\ \bibinfo
  {note} {[Erratum: Phys.Rev.Lett. 96, 039901 (2006)]},\ \Eprint
  {https://arxiv.org/abs/nucl-th/0410079} {arXiv:nucl-th/0410079} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Adamczyk}\ \emph {et~al.}(2017)\citenamefont
  {Adamczyk} \emph {et~al.}}]{STAR:2017ckg}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {L.}~\bibnamefont
  {Adamczyk}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Global $\Lambda$ hyperon polarization in nuclear
  collisions: evidence for the most vortical fluid}},\ }\href
  {https://doi.org/10.1038/nature23004} {\bibfield  {journal} {\bibinfo
  {journal} {Nature}\ }\textbf {\bibinfo {volume} {548}},\ \bibinfo {pages}
  {62} (\bibinfo {year} {2017})},\ \Eprint {https://arxiv.org/abs/1701.06657}
  {arXiv:1701.06657 [nucl-ex]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Adam}\ \emph {et~al.}(2018)\citenamefont {Adam} \emph
  {et~al.}}]{STAR:2018gyt}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.}~\bibnamefont
  {Adam}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Global polarization of $\Lambda$ hyperons in
  Au+Au collisions at $\sqrt{s_{_{NN}}}$ = 200 GeV}},\ }\href
  {https://doi.org/10.1103/PhysRevC.98.014910} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo {volume} {98}},\ \bibinfo
  {pages} {014910} (\bibinfo {year} {2018})},\ \Eprint
  {https://arxiv.org/abs/1805.04400} {arXiv:1805.04400 [nucl-ex]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Acharya}\ \emph {et~al.}(2020)\citenamefont {Acharya}
  \emph {et~al.}}]{ALICE:2019onw}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Acharya}} \emph {et~al.} (\bibinfo {collaboration} {ALICE}),\ }\bibfield
  {title} {\bibinfo {title} {{Global polarization of $\Lambda \bar \Lambda$
  hyperons in Pb-Pb collisions at $\sqrt {s_{NN}}$ = 2.76 and 5.02 TeV}},\
  }\href {https://doi.org/10.1103/PhysRevC.101.044611} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo {volume} {101}},\
  \bibinfo {pages} {044611} (\bibinfo {year} {2020})},\ \Eprint
  {https://arxiv.org/abs/1909.01281} {arXiv:1909.01281 [nucl-ex]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Abdallah}\ \emph {et~al.}(2021)\citenamefont
  {Abdallah} \emph {et~al.}}]{STAR:2021beb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.~S.}\ \bibnamefont
  {Abdallah}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Global $\Lambda$-hyperon polarization in Au+Au
  collisions at $\sqrt {s_{NN}}$=3~GeV}},\ }\href
  {https://doi.org/10.1103/PhysRevC.104.L061901} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo {volume} {104}},\
  \bibinfo {pages} {L061901} (\bibinfo {year} {2021})},\ \Eprint
  {https://arxiv.org/abs/2108.00044} {arXiv:2108.00044 [nucl-ex]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Abdulhamid}\ \emph {et~al.}(2023)\citenamefont
  {Abdulhamid} \emph {et~al.}}]{STAR:2023nvo}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.~I.}\ \bibnamefont
  {Abdulhamid}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Global polarization of \ensuremath{\Lambda} and
  \ensuremath{\Lambda}\textasciimacron{} hyperons in Au+Au collisions at
  sNN=19.6 and 27 GeV}},\ }\href {https://doi.org/10.1103/PhysRevC.108.014910}
  {\bibfield  {journal} {\bibinfo  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo
  {volume} {108}},\ \bibinfo {pages} {014910} (\bibinfo {year} {2023})},\
  \Eprint {https://arxiv.org/abs/2305.08705} {arXiv:2305.08705 [nucl-ex]}
  \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Liang}\ and\ \citenamefont
  {Wang}(2005{\natexlab{b}})}]{Liang:2004xn}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Z.-T.}\ \bibnamefont
  {Liang}}\ and\ \bibinfo {author} {\bibfnamefont {X.-N.}\ \bibnamefont
  {Wang}},\ }\bibfield  {title} {\bibinfo {title} {{Spin alignment of vector
  mesons in non-central A+A collisions}},\ }\href
  {https://doi.org/10.1016/j.physletb.2005.09.060} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume} {629}},\
  \bibinfo {pages} {20} (\bibinfo {year} {2005}{\natexlab{b}})},\ \Eprint
  {https://arxiv.org/abs/nucl-th/0411101} {arXiv:nucl-th/0411101} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Becattini}\ \emph {et~al.}(2013)\citenamefont
  {Becattini}, \citenamefont {Csernai},\ and\ \citenamefont
  {Wang}}]{Becattini:2013vja}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {F.}~\bibnamefont
  {Becattini}}, \bibinfo {author} {\bibfnamefont {L.}~\bibnamefont {Csernai}},\
  and\ \bibinfo {author} {\bibfnamefont {D.~J.}\ \bibnamefont {Wang}},\
  }\bibfield  {title} {\bibinfo {title} {{$\Lambda$ polarization in peripheral
  heavy ion collisions}},\ }\href {https://doi.org/10.1103/PhysRevC.88.034905}
  {\bibfield  {journal} {\bibinfo  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo
  {volume} {88}},\ \bibinfo {pages} {034905} (\bibinfo {year} {2013})},\
  \bibinfo {note} {[Erratum: Phys.Rev.C 93, 069901 (2016)]},\ \Eprint
  {https://arxiv.org/abs/1304.4427} {arXiv:1304.4427 [nucl-th]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Yang}\ \emph {et~al.}(2018)\citenamefont {Yang},
  \citenamefont {Fang}, \citenamefont {Wang},\ and\ \citenamefont
  {Wang}}]{Yang:2017sdk}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Y.-G.}\ \bibnamefont
  {Yang}}, \bibinfo {author} {\bibfnamefont {R.-H.}\ \bibnamefont {Fang}},
  \bibinfo {author} {\bibfnamefont {Q.}~\bibnamefont {Wang}},\ and\ \bibinfo
  {author} {\bibfnamefont {X.-N.}\ \bibnamefont {Wang}},\ }\bibfield  {title}
  {\bibinfo {title} {{Quark coalescence model for polarized vector mesons and
  baryons}},\ }\href {https://doi.org/10.1103/PhysRevC.97.034917} {\bibfield
  {journal} {\bibinfo  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo {volume}
  {97}},\ \bibinfo {pages} {034917} (\bibinfo {year} {2018})},\ \Eprint
  {https://arxiv.org/abs/1711.06008} {arXiv:1711.06008 [nucl-th]} \BibitemShut
  {NoStop}%
%%CITATION = ARXIV:1711.06008;%%
\bibitem [{\citenamefont {Sheng}\ \emph
  {et~al.}(2020{\natexlab{a}})\citenamefont {Sheng}, \citenamefont {Oliva},\
  and\ \citenamefont {Wang}}]{Sheng:2019kmk}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {X.-L.}\ \bibnamefont
  {Sheng}}, \bibinfo {author} {\bibfnamefont {L.}~\bibnamefont {Oliva}},\ and\
  \bibinfo {author} {\bibfnamefont {Q.}~\bibnamefont {Wang}},\ }\bibfield
  {title} {\bibinfo {title} {{What can we learn from the global spin alignment
  of $\phi$ mesons in heavy-ion collisions?}},\ }\href
  {https://doi.org/10.1103/PhysRevD.101.096005} {\bibfield  {journal} {\bibinfo
   {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {101}},\ \bibinfo
  {pages} {096005} (\bibinfo {year} {2020}{\natexlab{a}})},\ \bibinfo {note}
  {[Erratum: Phys.Rev.D 105, 099903 (2022)]},\ \Eprint
  {https://arxiv.org/abs/1910.13684} {arXiv:1910.13684 [nucl-th]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Xia}\ \emph {et~al.}(2021)\citenamefont {Xia},
  \citenamefont {Li}, \citenamefont {Huang},\ and\ \citenamefont
  {Zhong~Huang}}]{Xia:2020tyd}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {X.-L.}\ \bibnamefont
  {Xia}}, \bibinfo {author} {\bibfnamefont {H.}~\bibnamefont {Li}}, \bibinfo
  {author} {\bibfnamefont {X.-G.}\ \bibnamefont {Huang}},\ and\ \bibinfo
  {author} {\bibfnamefont {H.}~\bibnamefont {Zhong~Huang}},\ }\bibfield
  {title} {\bibinfo {title} {{Local spin alignment of vector mesons in
  relativistic heavy-ion collisions}},\ }\href
  {https://doi.org/10.1016/j.physletb.2021.136325} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Lett. B}\ }\textbf {\bibinfo {volume} {817}},\
  \bibinfo {pages} {136325} (\bibinfo {year} {2021})},\ \Eprint
  {https://arxiv.org/abs/2010.01474} {arXiv:2010.01474 [nucl-th]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Gao}(2021)}]{Gao:2021rom}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.-H.}\ \bibnamefont
  {Gao}},\ }\bibfield  {title} {\bibinfo {title} {{Helicity polarization in
  relativistic heavy ion collisions}},\ }\href
  {https://doi.org/10.1103/PhysRevD.104.076016} {\bibfield  {journal} {\bibinfo
   {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {104}},\ \bibinfo
  {pages} {076016} (\bibinfo {year} {2021})},\ \Eprint
  {https://arxiv.org/abs/2105.08293} {arXiv:2105.08293 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {M\"uller}\ and\ \citenamefont
  {Yang}(2022)}]{Muller:2021hpe}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {B.}~\bibnamefont
  {M\"uller}}\ and\ \bibinfo {author} {\bibfnamefont {D.-L.}\ \bibnamefont
  {Yang}},\ }\bibfield  {title} {\bibinfo {title} {{Anomalous spin polarization
  from turbulent color fields}},\ }\href
  {https://doi.org/10.1103/PhysRevD.105.L011901} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {105}},\
  \bibinfo {pages} {L011901} (\bibinfo {year} {2022})},\ \bibinfo {note}
  {[Erratum: Phys.Rev.D 106, 039904 (2022)]},\ \Eprint
  {https://arxiv.org/abs/2110.15630} {arXiv:2110.15630 [nucl-th]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Abdallah}\ \emph {et~al.}(2023)\citenamefont
  {Abdallah} \emph {et~al.}}]{STAR:2022fan}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M.~S.}\ \bibnamefont
  {Abdallah}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Pattern of global spin alignment of
  \ensuremath{\phi} and K$^{*0}$ mesons in heavy-ion collisions}},\ }\href
  {https://doi.org/10.1038/s41586-022-05557-5} {\bibfield  {journal} {\bibinfo
  {journal} {Nature}\ }\textbf {\bibinfo {volume} {614}},\ \bibinfo {pages}
  {244} (\bibinfo {year} {2023})},\ \Eprint {https://arxiv.org/abs/2204.02302}
  {arXiv:2204.02302 [hep-ph]} \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lv}\ \emph {et~al.}(2024)\citenamefont {Lv},
  \citenamefont {Yu}, \citenamefont {Liang}, \citenamefont {Wang},\ and\
  \citenamefont {Wang}}]{Lv:2024uev}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.-p.}\ \bibnamefont
  {Lv}}, \bibinfo {author} {\bibfnamefont {Z.-h.}\ \bibnamefont {Yu}}, \bibinfo
  {author} {\bibfnamefont {Z.-t.}\ \bibnamefont {Liang}}, \bibinfo {author}
  {\bibfnamefont {Q.}~\bibnamefont {Wang}},\ and\ \bibinfo {author}
  {\bibfnamefont {X.-N.}\ \bibnamefont {Wang}},\ }\bibfield  {title} {\bibinfo
  {title} {{Global quark spin correlations in relativistic heavy ion
  collisions}},\ }\href {https://doi.org/10.1103/PhysRevD.109.114003}
  {\bibfield  {journal} {\bibinfo  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo
  {volume} {109}},\ \bibinfo {pages} {114003} (\bibinfo {year} {2024})},\
  \Eprint {https://arxiv.org/abs/2402.13721} {arXiv:2402.13721 [hep-ph]}
  \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Sheng}\ \emph {et~al.}(2023)\citenamefont {Sheng},
  \citenamefont {Oliva}, \citenamefont {Liang}, \citenamefont {Wang},\ and\
  \citenamefont {Wang}}]{Sheng:2022wsy}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {X.-L.}\ \bibnamefont
  {Sheng}}, \bibinfo {author} {\bibfnamefont {L.}~\bibnamefont {Oliva}},
  \bibinfo {author} {\bibfnamefont {Z.-T.}\ \bibnamefont {Liang}}, \bibinfo
  {author} {\bibfnamefont {Q.}~\bibnamefont {Wang}},\ and\ \bibinfo {author}
  {\bibfnamefont {X.-N.}\ \bibnamefont {Wang}},\ }\bibfield  {title} {\bibinfo
  {title} {{Spin Alignment of Vector Mesons in Heavy-Ion Collisions}},\ }\href
  {https://doi.org/10.1103/PhysRevLett.131.042304} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {131}},\
  \bibinfo {pages} {042304} (\bibinfo {year} {2023})},\ \Eprint
  {https://arxiv.org/abs/2205.15689} {arXiv:2205.15689 [nucl-th]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Sheng}\ \emph
  {et~al.}(2020{\natexlab{b}})\citenamefont {Sheng}, \citenamefont {Wang},\
  and\ \citenamefont {Wang}}]{Sheng:2020ghv}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {X.-L.}\ \bibnamefont
  {Sheng}}, \bibinfo {author} {\bibfnamefont {Q.}~\bibnamefont {Wang}},\ and\
  \bibinfo {author} {\bibfnamefont {X.-N.}\ \bibnamefont {Wang}},\ }\bibfield
  {title} {\bibinfo {title} {{Improved quark coalescence model for spin
  alignment and polarization of hadrons}},\ }\href
  {https://doi.org/10.1103/PhysRevD.102.056013} {\bibfield  {journal} {\bibinfo
   {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {102}},\ \bibinfo
  {pages} {056013} (\bibinfo {year} {2020}{\natexlab{b}})},\ \Eprint
  {https://arxiv.org/abs/2007.05106} {arXiv:2007.05106 [nucl-th]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Sheng}\ \emph {et~al.}(2024)\citenamefont {Sheng},
  \citenamefont {Oliva}, \citenamefont {Liang}, \citenamefont {Wang},\ and\
  \citenamefont {Wang}}]{Sheng:2022ffb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {X.-L.}\ \bibnamefont
  {Sheng}}, \bibinfo {author} {\bibfnamefont {L.}~\bibnamefont {Oliva}},
  \bibinfo {author} {\bibfnamefont {Z.-T.}\ \bibnamefont {Liang}}, \bibinfo
  {author} {\bibfnamefont {Q.}~\bibnamefont {Wang}},\ and\ \bibinfo {author}
  {\bibfnamefont {X.-N.}\ \bibnamefont {Wang}},\ }\bibfield  {title} {\bibinfo
  {title} {{Relativistic spin dynamics for vector mesons}},\ }\href
  {https://doi.org/10.1103/PhysRevD.109.036004} {\bibfield  {journal} {\bibinfo
   {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {109}},\ \bibinfo
  {pages} {036004} (\bibinfo {year} {2024})},\ \Eprint
  {https://arxiv.org/abs/2206.05868} {arXiv:2206.05868 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Kumar}\ \emph {et~al.}(2023)\citenamefont {Kumar},
  \citenamefont {M\"uller},\ and\ \citenamefont {Yang}}]{Kumar:2023ghs}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.}~\bibnamefont
  {Kumar}}, \bibinfo {author} {\bibfnamefont {B.}~\bibnamefont {M\"uller}},\
  and\ \bibinfo {author} {\bibfnamefont {D.-L.}\ \bibnamefont {Yang}},\
  }\bibfield  {title} {\bibinfo {title} {{Spin alignment of vector mesons by
  glasma fields}},\ }\href {https://doi.org/10.1103/PhysRevD.108.016020}
  {\bibfield  {journal} {\bibinfo  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo
  {volume} {108}},\ \bibinfo {pages} {016020} (\bibinfo {year} {2023})},\
  \Eprint {https://arxiv.org/abs/2304.04181} {arXiv:2304.04181 [nucl-th]}
  \BibitemShut {NoStop}%
\bibitem [{\citenamefont {Chen}\ \emph {et~al.}(2017)\citenamefont {Chen},
  \citenamefont {Yang}, \citenamefont {Zhou},\ and\ \citenamefont
  {Liang}}]{Chen:2016iey}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {K.-b.}\ \bibnamefont
  {Chen}}, \bibinfo {author} {\bibfnamefont {W.-h.}\ \bibnamefont {Yang}},
  \bibinfo {author} {\bibfnamefont {Y.-j.}\ \bibnamefont {Zhou}},\ and\
  \bibinfo {author} {\bibfnamefont {Z.-t.}\ \bibnamefont {Liang}},\ }\bibfield
  {title} {\bibinfo {title} {{Energy dependence of hadron polarization in
  $e^+e^-\to hX$ at high energies}},\ }\href
  {https://doi.org/10.1103/PhysRevD.95.034009} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {95}},\ \bibinfo
  {pages} {034009} (\bibinfo {year} {2017})},\ \Eprint
  {https://arxiv.org/abs/1609.07001} {arXiv:1609.07001 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Li}\ \emph {et~al.}(2024)\citenamefont {Li},
  \citenamefont {Chen}, \citenamefont {Cao},\ and\ \citenamefont
  {Wei}}]{Li:2023qgj}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {X.}~\bibnamefont
  {Li}}, \bibinfo {author} {\bibfnamefont {Z.-X.}\ \bibnamefont {Chen}},
  \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Cao}},\ and\ \bibinfo
  {author} {\bibfnamefont {S.-Y.}\ \bibnamefont {Wei}},\ }\bibfield  {title}
  {\bibinfo {title} {{Correlations of dihadron polarization in central,
  peripheral, and ultraperipheral heavy-ion collisions}},\ }\href
  {https://doi.org/10.1103/PhysRevD.109.014035} {\bibfield  {journal} {\bibinfo
   {journal} {Phys. Rev. D}\ }\textbf {\bibinfo {volume} {109}},\ \bibinfo
  {pages} {014035} (\bibinfo {year} {2024})},\ \Eprint
  {https://arxiv.org/abs/2309.09487} {arXiv:2309.09487 [hep-ph]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Abelev}\ \emph {et~al.}(2007)\citenamefont {Abelev}
  \emph {et~al.}}]{STAR:2007ccu}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {B.~I.}\ \bibnamefont
  {Abelev}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Global polarization measurement in Au+Au
  collisions}},\ }\href {https://doi.org/10.1103/PhysRevC.76.024915} {\bibfield
   {journal} {\bibinfo  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo {volume}
  {76}},\ \bibinfo {pages} {024915} (\bibinfo {year} {2007})},\ \bibinfo {note}
  {[Erratum: Phys.Rev.C 95, 039906 (2017)]},\ \Eprint
  {https://arxiv.org/abs/0705.1691} {arXiv:0705.1691 [nucl-ex]} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Poskanzer}\ and\ \citenamefont
  {Voloshin}(1998)}]{Poskanzer:1998yz}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {A.~M.}\ \bibnamefont
  {Poskanzer}}\ and\ \bibinfo {author} {\bibfnamefont {S.~A.}\ \bibnamefont
  {Voloshin}},\ }\bibfield  {title} {\bibinfo {title} {{Methods for analyzing
  anisotropic flow in relativistic nuclear collisions}},\ }\href
  {https://doi.org/10.1103/PhysRevC.58.1671} {\bibfield  {journal} {\bibinfo
  {journal} {Phys. Rev. C}\ }\textbf {\bibinfo {volume} {58}},\ \bibinfo
  {pages} {1671} (\bibinfo {year} {1998})},\ \Eprint
  {https://arxiv.org/abs/nucl-ex/9805001} {arXiv:nucl-ex/9805001} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Adams}\ \emph {et~al.}(2003)\citenamefont {Adams}
  \emph {et~al.}}]{STAR:2003kbb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {J.}~\bibnamefont
  {Adams}} \emph {et~al.} (\bibinfo {collaboration} {STAR}),\ }\bibfield
  {title} {\bibinfo {title} {{Narrowing of the balance function with centrality
  in au + au collisions at (S(NN))**1/2 = 130-GeV}},\ }\href
  {https://doi.org/10.1103/PhysRevLett.90.172301} {\bibfield  {journal}
  {\bibinfo  {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {90}},\
  \bibinfo {pages} {172301} (\bibinfo {year} {2003})},\ \Eprint
  {https://arxiv.org/abs/nucl-ex/0301014} {arXiv:nucl-ex/0301014} \BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Bass}\ \emph {et~al.}(2000)\citenamefont {Bass},
  \citenamefont {Danielewicz},\ and\ \citenamefont {Pratt}}]{Bass:2000az}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.~A.}\ \bibnamefont
  {Bass}}, \bibinfo {author} {\bibfnamefont {P.}~\bibnamefont {Danielewicz}},\
  and\ \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Pratt}},\ }\bibfield
   {title} {\bibinfo {title} {{Clocking hadronization in relativistic heavy ion
  collisions with balance functions}},\ }\href
  {https://doi.org/10.1103/PhysRevLett.85.2689} {\bibfield  {journal} {\bibinfo
   {journal} {Phys. Rev. Lett.}\ }\textbf {\bibinfo {volume} {85}},\ \bibinfo
  {pages} {2689} (\bibinfo {year} {2000})},\ \Eprint
  {https://arxiv.org/abs/nucl-th/0005044} {arXiv:nucl-th/0005044} \BibitemShut
  {NoStop}%
\end{thebibliography}%

\documentclass[aps,superscriptaddress,showpacs]{revtex4-2}

\usepackage{graphicx}
\usepackage{epstopdf}
\usepackage{caption}
\usepackage{color}
\usepackage{bm}
\usepackage{textcomp} 
\usepackage{subfigure}
\usepackage{ulem}
\usepackage{color}
\usepackage{mathtools}
\usepackage{amsmath}
\graphicspath{{figures/}}

\begin{document}

\title{Considerations on measuring spin-spin correlation of hyeprons in heavy-ion experiments}

\author{Diyu Shen}
\email{dyshen@fudan.edu.cn}
\affiliation{Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China}
\author{Jinhui Chen}
\affiliation{Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China}
\author{Aihong Tang}
\affiliation{Brookhaven National Laboratory, Upton, New York 11973}
\date{July 2024}

\begin{abstract}
     The significant global spin alignment observed for $\phi$ mesons in heavy-ion collisions has sparked intense discussions about its origin and implications. One explanation suggests that fluctuations in the strong force field may introduce strong spin correlations between strange ($s$) and anti-strange ($\bar{s}$) quarks, leading to the global spin alignment of $\phi$ mesons. Extending this line of research, the theoretical community has proposed studying the spin correlation between $\Lambda$ and $\bar{\Lambda}$ hyperons. In this paper,  we construct experimental observables and make connections between them and the theoretical proposed quantities.
    
\end{abstract}
\maketitle
 
 The concept of global polarization in heavy-ion collisions was first proposed in Ref.\cite{Liang:2004ph}. In off-central heavy-ion collisions, quarks are predicted to be globally polarized along the direction of angular momentum due to spin-orbital coupling in strong interactions. This effect has recently been confirmed experimentally through measurements of $\Lambda$ global polarization~\cite{STAR:2017ckg,STAR:2018gyt,ALICE:2019onw,STAR:2021beb,STAR:2023nvo}. However, another related phenomenon, vector meson spin alignment, cannot be explained solely by vorticity or other  mechanisms, most of which are conventional~\cite{Liang:2004xn,Becattini:2013vja, Yang:2017sdk,Sheng:2019kmk,Xia:2020tyd,Gao:2021rom,Muller:2021hpe}. The measured $\rho_{00}$, the 00-component of the spin density matrix (with 1/3 indicating no spin alignment), of $\phi$ mesons is significantly above 1/3~\cite{STAR:2022fan}, while predictions from the quark coalescence model, considering $\Lambda$ data, suggest~\cite{Liang:2004xn, Yang:2017sdk}  $\rho_{00}$ below 1/3 with a much smaller deviation from 1/3. This discrepancy between hyperon polarization and vector meson spin alignment implies richer spin physics beyond global quark polarization. It was later realized that spin correlation between quarks and anti-quarks plays a crucial role in determining vector meson spin~\cite{Lv:2024uev}. For instance, short-range correlations between $s$ and $\bar{s}$ quarks due to strong-force fields successfully explain the observed $\phi$ meson spin alignment~\cite{Sheng:2022wsy,Sheng:2019kmk,Sheng:2020ghv,Sheng:2022ffb,Kumar:2023ghs}.

The spin-spin correlation of spin 1/2 particles was usually studied in helicity-frame in $e^+e^-$ collisions or $p+p$ collisons~\cite{Chen:2016iey,Li:2023qgj}, and the correlation function is defined as 
\begin{equation}
c_{ij} = \frac{f_{\uparrow\uparrow} + f_{\downarrow\downarrow} - f_{\downarrow\uparrow} - f_{\uparrow\downarrow}}{f_{\uparrow\uparrow} + f_{\downarrow\downarrow} + f_{\downarrow\uparrow} + f_{\uparrow\downarrow}},
\label{Eq:cnn}
\end{equation}
where $f_{m_im_j}=\langle m_im_j | \hat{\rho} | m_im_j \rangle$ denotes the fraction of fermion pairs in the spin state of $\left | m_im_j  \right \rangle$ along spin quantization direction, with $\hat{\rho}$ being the spin density matrix. In general, the quantization directions of particle $i$ and $j$ differ in helicity-frame. However, in heavy-ion collisions, we are interested in global spin-spin correlations, meaning the quantization directions of particle $i$ and $j$ are fixed. In this case, Eq.~\ref{Eq:cnn} is modified as~\cite{Lv:2024uev}:
\begin{equation}
c'_{ij} = c_{ij} - P_i P_j,
\label{Eq:cnn_prime}
\end{equation}
where $P_i$ and $P_j$ are the polarization of particle $i$ and $j$ respectively. It is clear that the $c'_{ij}$ represents $\langle \hat{\sigma}_{m_i} \hat{\sigma}_{m_j} \rangle - \langle \hat{\sigma}_{m_i} \rangle \langle \hat{\sigma}_{m_j} \rangle$, where $\hat{\sigma}_m$ is the Pauli operator along spin quantization direction. 

The left-hand side of Eq.~\ref{Eq:cnn_prime} represents the global spin correlations in heavy-ion collisions and can be calculated by measuring the $c_{ij}$ and global polarization. In the following, we demonstrate how to measure $c_{ij}$ using the example of $\Lambda$-$\Lambda$ spin correlation. This method can be easily extended to other hyperons. 

We begin with the proton distribution from $\Lambda$ weak decay,
\begin{equation}
    \frac{dN}{d\cos\theta^*}=\frac{1}{2}(1+\alpha_\Lambda P_\Lambda\cos\theta^*),
\end{equation}
where $\theta^*$ is the angle, in $\Lambda$'s rest frame, between proton momentum and $\Lambda$ spin quantization direction. $P_\Lambda$ defines $\Lambda$ polarization and $\alpha_\Lambda$ is a decay constant. For $\Lambda$ pairs in the spin state $\left | \uparrow\uparrow  \right \rangle$ (both are fully polarized thus $P_\Lambda=1$), the distribution of proton-pairs can be written as
\begin{equation}
\frac{dN_{\uparrow\uparrow}}{d\cos\theta_i^*d\cos\theta_j^*}=\frac{1}{4}\left [1+\alpha_\Lambda (\cos\theta_i^*+\cos\theta_j^*) + \alpha_\Lambda^2\cos\theta_i^*\cos\theta_j^* \right ],
\label{Eq:Npp}
\end{equation}
where $N_{\uparrow\uparrow}$ denotes proton pairs from $\Lambda$ pairs in $\left | \uparrow\uparrow  \right \rangle$ state. The footnote $i$ and $j$ denote the first proton and the second proton, respectively. Similarly, for $\Lambda$ pairs in the spin state of $\left | \downarrow\downarrow  \right \rangle$, $\left | \uparrow\downarrow  \right \rangle$, and $\left | \downarrow\uparrow  \right \rangle$, the proton pairs are described by 
\begin{eqnarray}
\frac{dN_{\downarrow\downarrow}}{d\cos\theta_i^*d\cos\theta_j^*}&=&\frac{1}{4}\left [1-\alpha_\Lambda (\cos\theta_i^*+\cos\theta_j^*) + \alpha_\Lambda^2\cos\theta_i^*\cos\theta_j^* \right ], \label{Eq:Nnn}\\
\frac{dN_{\uparrow\downarrow}}{d\cos\theta_i^*d\cos\theta_j^*}&=&\frac{1}{4}\left [1+\alpha_\Lambda (\cos\theta_i^*-\cos\theta_j^*) - \alpha_\Lambda^2\cos\theta_i^*\cos\theta_j^* \right ], \label{Eq:Npn}\\
\frac{dN_{\downarrow\uparrow}}{d\cos\theta_i^*d\cos\theta_j^*}&=&\frac{1}{4}\left [1-\alpha_\Lambda (\cos\theta_i^*-\cos\theta_j^*) - \alpha_\Lambda^2\cos\theta_i^*\cos\theta_j^* \right ]. \label{Eq:Nnp}
\end{eqnarray}
In reality, $\Lambda$ pairs are mixture of the above spin states, with the fractions given by $f_{m_im_j}$. Therefore, the general distribution of daughter proton pairs can be expressed as follows :
\begin{eqnarray}
\frac{dN}{d\cos\theta_i^*d\cos\theta_j^*} &=& f_{\uparrow\uparrow}\frac{dN_{\uparrow\uparrow}}{d\cos\theta_i^*d\cos\theta_j^*} + f_{\downarrow\downarrow}\frac{dN_{\downarrow\downarrow}}{d\cos\theta_i^*d\cos\theta_j^*} + f_{\uparrow\downarrow}\frac{dN_{\uparrow\downarrow}}{d\cos\theta_i^*d\cos\theta_j^*} + f_{\downarrow\uparrow}\frac{dN_{\downarrow\uparrow}}{d\cos\theta_i^*d\cos\theta_j^*} \notag \\
&=& \frac{1}{4}\left [ 1 + A\alpha_\Lambda\cos\theta_i^* 
+ B\alpha_\Lambda\cos\theta_j^* +C\alpha_\Lambda^2\cos\theta_i^*\cos\theta_j^* \right ],
\label{Eq:theta}
\end{eqnarray}
where 
\begin{eqnarray}
A=\frac{f_{\uparrow\uparrow}-f_{\downarrow\downarrow}+f_{\uparrow\downarrow}-f_{\downarrow\uparrow}}{f_{\uparrow\uparrow} + f_{\downarrow\downarrow} + f_{\downarrow\uparrow} + f_{\uparrow\downarrow}}, \notag \\
B=\frac{f_{\uparrow\uparrow}-f_{\downarrow\downarrow}-f_{\uparrow\downarrow}+f_{\downarrow\uparrow}}{f_{\uparrow\uparrow} + f_{\downarrow\downarrow} + f_{\downarrow\uparrow} + f_{\uparrow\downarrow}}, \notag \\
C=\frac{f_{\uparrow\uparrow}+f_{\downarrow\downarrow}-f_{\uparrow\downarrow}-f_{\downarrow\uparrow}}{f_{\uparrow\uparrow} + f_{\downarrow\downarrow} + f_{\downarrow\uparrow} + f_{\uparrow\downarrow}}.
\end{eqnarray}
The parameter $C$ is the $c_{ij}$ defined in Eq.~\ref{Eq:cnn}. Indeed, the parameter $A$ could be reduced to $(f_{\uparrow}-f_{\downarrow})/(f_{\uparrow}+f_{\downarrow})$ which represents the polarization of the first $\Lambda$ by definition, where $f_{\uparrow}=f_{\uparrow\uparrow}+f_{\uparrow\downarrow}$ ($f_{\downarrow}=f_{\downarrow\uparrow}+f_{\downarrow\downarrow}$) is the fraction of it in the spin state $\left | \uparrow  \right \rangle$ ($\left | \downarrow  \right \rangle$). Similarly, the parameter $B$ represents the polarization of the second $\Lambda$. It is reasonable to expect that parameter $C$ reflects the spin correlation between the first and second $\Lambda$. 
Mathematically, the $c_{ij}$ can be obtained by the integral of $\cos\theta^*_i\cos\theta^*_j$, it is equivalent to calculate the average $\langle \cos\theta^*_i\cos\theta^*_j \rangle$ in the experiment. 

Based on Eq.~\ref{Eq:cnn_prime}, we establish the connection between experimental observable and the $\Lambda$ spin correlation as used in the context of theoretical work: 
\begin{equation}
c'_{\Lambda\Lambda}=\frac{9}{\alpha^2_\Lambda}\langle \cos\theta^*_i\cos\theta^*_j \rangle - P_\Lambda^2.
\end{equation}
In experiment, the spin quantization direction (orbital angular momentum direction) is estimated from the event plane, which has a finite resolution. To correct for event plane resolution, similar to Ref.~\cite{STAR:2007ccu}, we project Eq.~\ref{Eq:theta} onto transverse plane, resulting in :
\begin{equation}
\frac{dN}{d\Delta\phi_i^*d\Delta \phi_j^*} = \frac{1}{16\pi}(1+A\alpha_\Lambda\sin\Delta \phi_i^* + B\alpha_\Lambda\sin\Delta \phi_j^*+\frac{C\alpha_\Lambda^2\pi}{4}\sin\Delta \phi_i^*\sin\Delta \phi_j^*),
\end{equation}
and $c'_{\Lambda\Lambda}$ can be calculated by
\begin{equation}
c'_{\Lambda\Lambda}=\frac{64}{\pi^2\alpha_\Lambda^2}\langle\sin\Delta\phi_i^*\sin\Delta \phi_j^* \rangle - P_\Lambda^2,
\label{Eq:FreeStyle}
\end{equation}
where $\Delta \phi_i^*$ ($\Delta \phi_j^*$) is the azimuthual angle relative to the reaction plane for proton $i$ ($j$) in first (second) $\Lambda$ rest frame. Considering a finite event plane resolution, Eq.~\ref{Eq:FreeStyle} can be expressed as
\begin{equation}
  c'_{\Lambda\Lambda}=\frac{64}{\pi^2\alpha_\Lambda^2}\frac{\langle\sin(\phi_i^*-\Psi_{\rm EP})\sin(\phi_j^*-\Psi_{\rm EP}) \rangle}{\langle \cos^2(\Psi_{\rm EP} - \Psi_{\rm RP}) \rangle} - P_\Lambda^2,  
\end{equation}
where $\Psi_{\rm EP}$ and $\Psi_{\rm RP}$ denote the event plane and the reaction plane, respectively. Note that
\begin{equation}
   \langle \cos^2(\Psi_{\rm EP} - \Psi_{\rm RP}) \rangle =   \frac{1 + \langle \cos(2(\Psi_{\rm EP} - \Psi_{\rm RP})) \rangle }{2},
\end{equation}
and $\langle \cos(2(\Psi_{\rm EP} - \Psi_{\rm RP})) \rangle$ can be determined from sub-event method in experiment~\cite{Poskanzer:1998yz}.

Building on the structure of balance function~\cite{STAR:2003kbb,Bass:2000az}, we propose a spin balance function to detect the difference of spin correlation between ``same sign'' pairs ($\Lambda$-$\Lambda$ or $\bar{\Lambda}$-$\bar{\Lambda}$) and ``opposite sign'' pairs ($\Lambda$-$\bar{\Lambda}$),
\begin{equation}
B(c'_{OS},c'_{SS}) = \frac{1}{2}(\frac{c'_{\Lambda\bar{\Lambda}} - c'_{\Lambda\Lambda}}{c'_{\Lambda\bar{\Lambda}} + c'_{\Lambda\Lambda}} + \frac{c'_{\bar{\Lambda}\Lambda} - c'_{\bar{\Lambda}\bar{\Lambda}}}{c'_{\bar{\Lambda}\Lambda} + c'_{\bar{\Lambda}\bar{\Lambda}}} ).
\end{equation}
The advantage of the above equation is that the detector effect and flavor independent correlation will be canceled, only additional correlation between $s$ and $\bar{s}$ (in $\Lambda$-$\bar{\Lambda}$ pairs) remains.

To summarize, in this paper, we investigated the correlation function used in a theoretical context, identified its connection to experimental observables, and outlined the procedure to measure it in experiments.  Additionally, we proposed a method for measuring  the  correlation using a spin balance function, which minimize the impact of detector and other trivial effects. This work will be instrumental in studying hyperon spin correlations and facilitating the understanding of spin-spin correlations and their implications for fluctuations in the strong force field. \\

\section*{Acknowledgment}
This work was inspired by the theoretical study \cite{Lv:2024uev} and benefited from discussions with X.-N. Wang, Z.-T. Liang, Q. Wang, X.-L. Sheng., and Z. Tu.

\bibliography{ref}

\end{document}